OpenStudy (anonymous):

Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x^2 + 7x + 3 = 0 Numerical Answers Expected!

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

I think I might suicide.

Parth (parthkohli):

You should expect numerical answers from yourself though

OpenStudy (anonymous):

if \[b^2-4ac\] is a perfect square, then your solutions are rational numbers so you can factor over the integers

OpenStudy (anonymous):

Well that makes no sense.

Parth (parthkohli):

The discriminant

OpenStudy (anonymous):

to me

Parth (parthkohli):

Do you know what discriminant is?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

actually it makes perfect sense the solutions to \[ax^2+bx+c=0\] are \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

Parth (parthkohli):

^ that's the quadratic formula @careless850

OpenStudy (anonymous):

Yeah, I knew that.

Parth (parthkohli):

The number of solutions depend on the discriminant.

OpenStudy (anonymous):

I don't know what part of that formula to look at to determine when to factor though.

OpenStudy (anonymous):

inside the radical sits the expression \[b^2-4ac\] and if that expression is not a perfect square, like say 9 or 16, then your solutions are irrational numbers because \[\sqrt{\text{whatever}}\] is only rational if "whatever" is a perfect square

Parth (parthkohli):

If the discriminant is 0, then it has one solution. If the discriminant is more than 0, then it has real solutions. If the discriminant is less than 0, then it has complex solutions.

OpenStudy (anonymous):

And a perfect square is two numbers that both sit on the axis, right?

Parth (parthkohli):

^ if I correctly remember

OpenStudy (anonymous):

so if you compute \[b^2-4ac\] and get for example the number 5, then the two solutions are irrational and you will not be able to factor using integers

Parth (parthkohli):

What are b, a and c here? @careless850

OpenStudy (anonymous):

no a perfect square is not a number that sits on the real axis. a perfect square is a number like \[0,1,4,9,16,25,36,49,64,...\]

OpenStudy (anonymous):

2 = a. 7 = b. and 3 = c.

Parth (parthkohli):

Then compute \(b^2 - 4ac\)

OpenStudy (anonymous):

7*7 = 49.

Parth (parthkohli):

\( \color{Black}{\Rightarrow (7)^2 - 4(2)(3) }\)

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

49-24.

OpenStudy (anonymous):

25.

Parth (parthkohli):

And that's a perfect square :)

OpenStudy (anonymous):

Because 5 * 5 is 25, right?

Parth (parthkohli):

Yes, exactly! :D

OpenStudy (anonymous):

or 5^2?

Parth (parthkohli):

5^2 = 5 * 5

OpenStudy (anonymous):

Alright, but what does this have to do with what my problem is asking...?

Parth (parthkohli):

You got your answer.

Parth (parthkohli):

1) b^2 - 4ac as sat already told you.

Parth (parthkohli):

2) 25

OpenStudy (anonymous):

Alright.